They did not give any information about it they just ask duration with those informationeste s temmepw Calculdbons Dogukan Kara DK Transitions Animations Slide Show Review View Help Search Share og Layout - E Rese
Answer #1

Solution is being provided asssuming a face value of 1,000.

CF for Year 1,2,3=Face avlue*3%=30

CF for year 4=Redemption value+coupon=1000+30=1030

Discount factor=1/(1+YTM)n where n is number of year.

Price=Sum total of PV column=929.07

Duration=3.82 years (as shown in table below).

Year CF Discount factor PV PV/Price (PV/Price)*T
1 30 0.9523 28.57 0.03075 0.03075
2 30 0.9070 27.21 0.02929 0.05858
3 30 0.8638 25.91 0.02789 0.08366
4 1030 0.8227 847.38 0.9121 3.6483
TOTAL 929.07 3.8212